OpenStudy (anonymous):

Find the value of the square root of 8 over 3 times the square root of 4. in simplest form. 6 2 over 3. the square root of 2 over 3. 3 times the square root of 2.

OpenStudy (anonymous):

\[\frac{\sqrt{8}}{3}\sqrt{4}=\frac{\sqrt{32}}{3}=\frac{\sqrt{16*2}}{3}=\frac{4\sqrt{2}}{3}\]

OpenStudy (anonymous):

i think they mean \[\frac{\sqrt 8}{3\sqrt4}\]

OpenStudy (anonymous):

Hard to read. Will correct :)

OpenStudy (anonymous):

also, it is better to explain what you did than to give the answer

OpenStudy (anonymous):

\[\frac{\sqrt{8}}{3\sqrt{4}}\] Times by sqrt 4 on top and bottom: \[\frac{\sqrt{32}}{3*4}=\frac{\sqrt{16*2}}{12}=\frac{4\sqrt{2}}{12}\] Simplify by dividing numerator and denominator by 4: \[\frac{\sqrt{2}}{3}\]

OpenStudy (anonymous):

note that \[\sqrt 8 = \sqrt{4 \times 2} = \sqrt 4 \times \sqrt 2\]

OpenStudy (anonymous):

Not to be picky but the steps were there. If I skipped straight to the answer you'd see just an answer like other posts.

OpenStudy (anonymous):

I agree that it could be clarified more but people are welcome to ask questions of explanations given.

OpenStudy (anonymous):

yes thats true :)

OpenStudy (anonymous):

√2 = a/b so 2 = a2/b2 so 2b2 = a2 This means that a2 is even, and so a must be even. So we can replace a by 2c. so 2b2 = (2c)2 so 2b2 = 4c2 so b2 = 2c2

OpenStudy (anonymous):

is it write

OpenStudy (anonymous):

andremon